# MedMax: Mixed-Modal Instruction Tuning for Training Biomedical Assistants

### v1

@@ Acknowledgements
Appendix A Limitations
@@
Furthermore, we emphasize that the MedMax model is a research prototype designed to encourage community engagement in building capable biomedical assistants. In cases of medical emergencies, we strongly advise users to consult a healthcare professional rather than relying solely on the model’s output. With ongoing community feedback and the collection of high-quality expert data, we aim to improve the model’s faithfulness and safety over time.
Appendix B MedMax-Instruct Data Curation Prompts
@@ Appendix F Additional VQA Results
Appendix G Finetuning Details
@@ Appendix H Additional Image Captioning Results
We compare the performance of the MedMax model with other relevant captioning models on the image captioning task using BioCLIPMedScore. The results are presented in Table 15. Our analysis reveals that the MedMax model outperforms LLaVA Med-v1.5 by an average of $0.16$ points. Furthermore, we find that the performance of the MedMax model is on par with diverse baselines such as HuatuoGPT-Vision and GPT-4o for the captioning task. Overall, this underscores the capability of the MedMax model in training robust mixed-modal models that excel in diverse biomedical tasks.

### v2

@@ Acknowledgements
Appendix A More Details on Data Curation
A.1 LLaVA-Med-PMC
A.2 PMC-OA
Further, we utilize the PMC-OA data that also consists of image-caption data curated from PubMedCentral research papers. Its notable features includes its accessibility (open-access), size ($1$M+ instances) and diversity of biomedical imaging data (e.g., ultrasound, fMRI, endoscope, PET). To maintain one-to-one correspondence in the data, we filter instances where a caption was aligned with multiple sub-images. Further, we filter small biomedical images that were less than 200 pixels in width or height in this data.
A.3 PubMedVision
In many cases, the real image-caption has inherent data noise and formatting issues. Hence, we include the synthetically-generated PubMedVision-AlignmentVQA [64] data which utilizes GPT-4-Vision [38] to denoise and reformat noisy internet data for biomedicine. In particular, we filter the original data ($647$K) to remove multiple image instances in this data to get $504$K instances, and randomly select a subset of $100$K instances for the MedMax mix. While this dataset can be utilized for image captioning (input image and question as context, and the answer as context), it could not be utilized for image generation directly. To address this, we prompt GPT-4o-mini [16] to convert the image descriptions in the PubMedVision-AlignmentVQA data into image generation prompts: ‘Convert the image description into an image generation prompt with AI’. We show an example in Appendix Table 6.
A.4 MIMIC-CXR
We collect chest radiographs along with the medical reports in the MIMIC-CXR [18] data. Originally, the dataset consisted of 377K instances. We filtered it to exclude reports discussing more than one image, reducing the dataset to 102K instances. Subsequently, we subsampled the data to decrease the proportion of ‘No findings’ reports [7] from $20\%$ to $10\%$.
Appendix B Limitations
@@
Appendix C MedMax-Instruct Data Curation Prompts
@@ Appendix F Additional VQA Results
Appendix G Model and Finetuning Details
Chameleon [34] represents the raw images as discrete visual tokens using a VQGAN [11], and the text data into discrete text tokens using BPE tokenizer [44]. Subsequently, each instance in the training dataset is represented as a sequence of discrete tokens, and the model is trained to predict the next token based on the preceding tokens in the sequence, following an autoregressive objective. The model consists of a vocabulary size of $65536$ where $8192$ are visual tokens apart from beginning of image and end of image tokens. In addition, the vocabulary includes a reserved token ‘<reserved08706>’ that separates the instruction (context) from the response (output) for instruction-tuning. Post-tokenization, the entire MedMax data consists $1.7$B tokens where $0.7$B and $1$B are visual and text tokens, respectively.
While the Chameleon-7B model weights are publicly available, they were safety-tuned and support mixed-modal inputs and text-only output to be used for research purposes. [^4] https://ai.meta.com/blog/meta-fair-research-new-releases/ To unlock the mixed-modal output capabilities, Anole [9] selectively finetunes the output embeddings of the image tokens using high-quality images from LAION [43]. This strategy does not interfere with the input mixed-modal understanding and text-only output abilities of the original Chameleon model.
@@ Appendix H Additional Image Captioning Results
We compare the performance of the MedMax model with other relevant captioning models on the image captioning task using BioCLIPMedScore. The results are presented in Table 15. We observe that the average BioMedCLIPScore for our model is better LLaVA Med-v1.5 and at par with baselines such as HuatuoGPT-Vision, GPT-4o-mini, and GPT-4o. Overall, this underscores the capability of the MedMax model in training robust mixed-modal models that excel in diverse biomedical tasks.
@@
Appendix I Qualitative Examples
We present the qualitative examples in Figure 12 and Figure 13.
Figure 12: Positive qualitative examples. Sample generations from the MedMax model across diverse tasks: multimodal (interleaved) generation, visual chat, image captioning, and image generation. We verify with experts cited in acknowledgments that the provided medical information is high quality and correct.
Figure 13: Negative qualitative examples. Sample generations from the MedMax model across diverse tasks: multimodal (interleaved) generation, visual chat, image captioning, and image generation. Pitfalls include poor image generation, confusion between segments within an image, and misaligned captions.